Three capacitor of capacities 2*10-6F,4*10-6F and 8*10-3 are joined in series and the combination is connected to a 12 V source.Calculate the energy stored in the capacitor when thse are fully charged.

Grade:12

1 Answers

Vikas TU
5 years ago
The resulatant capicatoance would be =>
1/C  = 1/C1 + 1/C2 + 1/C3
1/C = 1/(2*10^-6) +  1/(4*10^-6) + 1/(8*10^-3) => (0.5 + 0.25 + 0.125)10^6 =>0.875*10^6
Ceq. = 1.142 x 10^-6 F
 
Energy storred = 0.5CV^2 => 0.5*1.142 x 10^-6*144 => 82.244*10^-6 J